# Two-Level Distributed Interference Management for Large-Scale HAPS-Empowered vHetNets

### v1

@@ IV-B1 Inner-level ADMM block 1 optimization problem
In this way, the right hand side of inequality (12a) will be replaced with by $p_{u}^{2}+q_{u}^{2}$. Accordingly, moving $\beta_{u}$ to the right hand side and replacing the right hand side with its approximated first-order Taylor series, the constraint can be approximated as
@@
Algorithm 1 describes the first-block SCA-based algorithm. Algorithm 1 solves the approximated problem (IV-B1) iteratively for the optimal solution, which is the suboptimal solution of the original problem (29b). The algorithm stops upon convergence, which is defined as when a tolerable change in the objective function value is achieved (less than $\epsilon$), or the maximum number of SCA iterations, $N_{\text{iter}}$, is reached, whichever first. This algorithm can be solved centrally or distributedly at each BS using the same initial and broadcasted values.
Algorithm 1 SCA-based algorithm for ADMM Block 1.
1:  Initialize $\mathbf{p}^{(0)}$, $\mathbf{q}^{(0)}$, and $\boldsymbol{\beta}^{(0)},$ set $m\leftarrow 0$.
@@
3:     Solve (IV-B1) to find $\mathcal{X}^{(m)*}$, $\mathbf{p}^{(m)*}$, $\mathbf{q}^{(m)*}$.
4:     Update $\mathbf{p}^{(m+1)}\leftarrow\mathbf{p}^{(m)*}$, $\mathbf{q}^{(m+1)}\leftarrow\mathbf{q}^{(m)*}$, $m\leftarrow m+1$.
@@ IV-B2 Inner-level ADMM block 2 optimization problem
$\overline{I}_{b,u}\geq\sum_{\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|}^{2},~{}\forall u,$
@@ V-A1 Stopping Criteria for the inner-level algorithm 2
00 0 $\in$ $\nabla f(\mathcal{X}^{(t)})+\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t-1)}_{l,b}+\sum_{b\in\mathcal{B}}\Bigl{[}\rho^{(k)}_{s,b}\bigl{(}\mathbf{s}^{(t)}_{b}-\overline{\mathbf{s}}^{(t-1)}_{b}+\mathbf{z}^{(t-1)}_{s,b}\bigr{)}+\rho^{(k)}_{I,b}\bigl{(}\mathbf{I}^{(t)}_{b}-\overline{\mathbf{I}}^{(t-1)}_{b}+\mathbf{z}^{(t-1)}_{I,b}\bigr{)}\Bigr{]}+N_{\mathcal{C}}(\mathcal{X}^{(t)}),\qquad$ (37a)
00 0 $\in$ $-\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t-1)}_{l,b}-\sum_{b\in\mathcal{B}}\Bigl{[}\rho^{(k)}_{s,b}\bigl{(}\mathbf{s}^{(t)}_{b}-\overline{\mathbf{s}}^{(t)}_{b}+\mathbf{z}^{(t-1)}_{s,b}\bigr{)}+\rho^{(k)}_{I,b}\bigl{(}\mathbf{I}^{(t)}_{b}-\overline{\mathbf{I}}^{(t)}_{b}+\mathbf{z}^{(t-1)}_{I,b}\bigr{)}\Bigr{]}+N_{\overline{\mathcal{C}}}(\overline{\mathcal{X}}^{(t)}).$ (37b)
@@
$\sum_{b\in\mathcal{B}}\Bigl{[}\rho^{(k)}_{s,b}\bigl{(}-\overline{\mathbf{s}}^{(t)}_{b}+\mathbf{z}^{(t)}_{s,b}+\overline{\mathbf{s}}^{(t-1)}_{b}-\mathbf{z}^{(t-1)}_{s,b}\bigr{)}+\rho^{(k)}_{I,b}\bigl{(}-\overline{\mathbf{I}}^{(t)}_{b}+\mathbf{z}^{(t)}_{I,b}+\overline{\mathbf{I}}^{(t-1)}_{b}-\mathbf{z}^{(t-1)}_{I,b}\bigr{)}\Bigr{]}\in\nabla f(\mathcal{X}^{(t)})+\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t)}_{l,b}+N_{\mathcal{C}}(\mathcal{X}^{(t)}),\qquad$ (38a)
@@
Comparing (38a) and (36a), the stopping criteria for inner-level algorithm 2 will be according to (35a).
@@ V-B Convergence
In order to discuss the convergence of the proposed algorithm, we make below three assumptions, which hold for the formulated problem (IV-A), considering sufficient available transmit power at each BS.
@@
Assumption 2.
@@ V-C Computational Complexity and Signaling Overhead
VI Numerical Result and Discussion
@@
HAPS altitude $20$ Km
Center frequency ($f_{c}$) $2.545$ GHz
@@
$d_{x},~{}d_{y}$ $\lambda/2$
$\sigma^{2}_{n}$ $-100$ dBm
$P_{b}^{\text{max}},~{}\forall b\in\mathcal{B}\setminus\{B+1\},~{}P_{B+1}^{\text{max}}$ $43$ dBm, $52$ dBm
Outer-level penalty parameter ($\rho^{(0)}_{o}$) $10$
$\delta$ $2$
$\omega$ $0.5$
$\gamma$ $1.5$
@@
In this section, we evaluate the performance of the proposed distributed PFBWD algorithm in HAPS-empowered vHetNets and compare it against both standalone terrestrial networks and the centralized approach. The standalone terrestrial network refers to a scenario where no HAPS is involved. In the case of a centralized algorithm, the original problem is solved using the SCA framework, which incorporates some RLT, as detailed in [4]. The simulated network consists of four MBSs and one HAPS, deployed over a $4~{}\text{km}\times 4~{}\text{km}$ square urban area. The HAPS and MBSs are equipped with $8\times 8$ and $4\times 4$ MIMO antenna arrays, respectively, and jointly serve $U$ single-antenna UEs in a cell-free architecture. The UEs are uniformly distributed across the considered area. The remaining simulation parameters and their values are summarized in Table I. All results are averaged over $1000$ independent network realizations. The algorithm is implemented using CVX parser, which runs on MATLAB, with mosek 9.1.9 as an internal solver [38].
@@
(a) CDF of average SE per UE (b/s/Hz/UE).
(b) CDF of average PF objective function value.
Figure 4: Statistical behavior of SE per UE and the average PF objective function value in a HAPS-empowered vHetNet ($4$ MBSs + $1$ HAPS) serving $16$ UEs.
(a) Average SE per UE vs. number of UEs
(b) Average PF objective function value vs. number of UEs
Figure 5: Performance comparison of centralized and distributed PFBWD algorithms for different numbers of UEs in a HAPS-empowered vHetNet ($4$ MBSs + $1$ HAPS).
@@
First, we discuss the convergence behavior of the proposed distributed PFBWD algorithm. Fig. 3 plots the average sum of residuals, defined as $\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\left\|\mathbf{r}_{l,b}\right\|$, over the inner-level ADMM Algorithm 2 iterations $t$, and the average PF objective function value (i.e., $\sum_{u\in\mathcal{U}}{\log(\log_{2}(1+\gamma_{u}))}$) over the outer-level Algorithm 3 iterations $k$. It can be observed that the residual values converge to zero as $t$ increases, while the objective value increases with $k$, demonstrating the convergence behavior of the proposed algorithm. The outer-level ALM algorithm also exhibits fast convergence as the objective value saturates after a few iterations (on an average $k=4$ iterations). However, the choice of penalty parameters plays a critical role in determining the number of required iterations. As shown in Table II, as the ratio of the inner-level to outer-level penalty parameters (i.e., $\delta$ in Algorithm 2) increases the number of required iterations reduces. This is because a higher penalty parameter enforces the satisfaction of the coupling constraints more aggressively, thereby accelerating convergence to feasibility. However, excessively large values of $\delta$ may lead to numerical instability [24].
Figs. 4a and  4b plot the cumulative distribution function (CDF) of the average SE per UE and the average PF objective value, respectively. The results compare the proposed distributed PFBWD algorithm with the centralized algorithm in a vHetNet comprising $4$ MBSs and $1$ HAPS, serving $16$ UEs. It can be observed that the distributed algorithm achieves lower average SE and objective function values. This is because the centralized algorithm has access to all network information, including complete channel coefficient matrices, enabling it to make decisions closer to the optimal solution. Conversely, in the distributed algorithm, each BS makes local decisions based on its channel information and coordinates with others by sharing selected variables. Consequently, the distributed approach yields slightly lower performance than the centralized counterpart. However, this performance gap is small, and the distributed algorithm’s results remain comparable to the centralized solution.
@@
TABLE III: Comparison of computational complexity and signaling overhead for centralized vs. distributed PFBWD algorithms in a vHetNet ($4$ MBSs + $1$ HAPS) serving $16$ UEs.
@@
Furthermore, Fig. 5, illustrates the performance gap between the centralized and distributed PFBWD algorithms with increasing number of UEs. Specifically, Figs. 5(a) and  5(b) plot the average SE per UE and the average PF objective function value, respectively, in a HAPS-empowered vHetNet consisting of $4$ MBSs and $1$ HAPS. As the number of UEs grows, the average SE per UE decreases mainly due to two reasons: i) the available resources are shared among more UEs and ii) the increased interference. Additionally, the performance gap between the centralized and distributed algorithms slightly widens with increasing UEs. This is because the amount of missing information at each BS in the distributed approach grows, leading to a minor decline in performance. Importantly, this minor decrease in performance comes with significant gains in algorithmic complexity and reduced signaling overhead among BSs. Table III summarizes the computational complexity and signaling overhead requirements for both algorithms when serving $16$ UEs, considering $4\times 4$ and $8\times 8$ antenna arrays at MBSs and HAPS, respectively. The proposed PFBWD Algorithm 3 offers much lower complexity and substantially less data exchange, resulting in a significant reduction in signaling overhead.
@@
Figure 7: Impact of HAPS antenna architecture on the average SE per UE in a vHetNet ($4$ MBSs + $1$ HAPS) serving $16$ UEs.
In line with this observation, Fig. 7 presents the average SE per UE in a $(4+1)$ vHetNet configuration serving $16$ UEs, for various HAPS antenna configurations. As shown, the SE gain provided by HAPS integration can be further amplified by increasing the number of antenna elements at the HAPS. This is particularly promising, as the large surface area of HAPS platforms allows for accommodating a significantly larger number of antennas, enabling improved beamforming capabilities and higher spatial resolution.

### v2

@@ IV-B1 Inner-level ADMM block 1 optimization problem
In this way, the right-hand side of inequality (14a) will be replaced with $p_{u}^{2}+q_{u}^{2}$. Accordingly, moving $\beta_{u}$ to the right-hand side and replacing the right-hand side with its approximated first-order Taylor series, the constraint can be approximated as
@@
Algorithm 1 describes the first-block SCA-based algorithm. At each inner-level ADMM iteration $t$, Algorithm 1 solves the approximated problem (IV-B1) iteratively for the optimal solution. The optimal point of the SCA-based approximated problem satisfies the Karush-Kuhn-Tucker (KKT) conditions of the original problem (30b) [[30] Proposition 3.2]. Therefore, the solution of Algorithm 1 is a stationary point of the problem (30b), and hence a feasible solution. The SCA-based Algorithm 1 stops upon convergence, which is defined as when a tolerable change in the objective function value is achieved (less than $\epsilon$), or the maximum number of SCA iterations, $N_{\text{iter}}$, is reached, whichever first. This algorithm can be solved centrally or in a distributed manner at each BS using the same initial and broadcasted values.
Algorithm 1 SCA-based algorithm for ADMM Block 1 iteration $t$.
1: Initialize $\mathbf{p}^{(0)}$, $\mathbf{q}^{(0)}$, $\boldsymbol{\beta}^{(0)},$ and $m\leftarrow 0$.
@@
3:  Solve (IV-B1) to find $\mathcal{X}^{(t)*}$.
4:  Update $\mathbf{p}^{(m+1)}\leftarrow\mathbf{p}^{(m)*}$, $\mathbf{q}^{(m+1)}\leftarrow\mathbf{q}^{(m)*}$, $\boldsymbol{\beta}^{(m+1)}\leftarrow\boldsymbol{\beta}^{(m)^{*}},$ $m\leftarrow m+1$.
@@ IV-B2 Inner-level ADMM block 2 optimization problem
$\overline{I}_{b,u}\geq\sum_{k\in\mathcal{U}\setminus\{u\}}{\left|\left(\mathbf{h}^{b}_{u}\right)^{H}\mathbf{w}^{b}_{k}\right|}^{2},~\forall u,$
@@ V-A1 Stopping Criteria for the inner-level algorithm 2
$0\in\nabla f(\mathcal{X}^{(t)})+\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t-1)}_{l,b}+\sum_{b\in\mathcal{B}}\Bigl[\rho^{(k)}_{s,b}\bigl(\mathbf{s}^{(t)}_{b}-\overline{\mathbf{s}}^{(t-1)}_{b}+\mathbf{z}^{(t-1)}_{s,b}\bigr)+\rho^{(k)}_{I,b}\bigl(\mathbf{I}^{(t)}_{b}-\overline{\mathbf{I}}^{(t-1)}_{b}+\mathbf{z}^{(t-1)}_{I,b}\bigr)\Bigr]+N_{\mathcal{C}}(\mathcal{X}^{(t)}),\qquad$ (38a)
$0\in-\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t-1)}_{l,b}-\sum_{b\in\mathcal{B}}\Bigl[\rho^{(k)}_{s,b}\bigl(\mathbf{s}^{(t)}_{b}-\overline{\mathbf{s}}^{(t)}_{b}+\mathbf{z}^{(t-1)}_{s,b}\bigr)+\rho^{(k)}_{I,b}\bigl(\mathbf{I}^{(t)}_{b}-\overline{\mathbf{I}}^{(t)}_{b}+\mathbf{z}^{(t-1)}_{I,b}\bigr)\Bigr]+N_{\overline{\mathcal{C}}}(\overline{\mathcal{X}}^{(t)}).$ (38b)
@@
$\sum_{b\in\mathcal{B}}\Bigl[\rho^{(k)}_{s,b}\bigl(-\overline{\mathbf{s}}^{(t)}_{b}+\mathbf{z}^{(t)}_{s,b}+\overline{\mathbf{s}}^{(t-1)}_{b}-\mathbf{z}^{(t-1)}_{s,b}\bigr)+\rho^{(k)}_{I,b}\bigl(-\overline{\mathbf{I}}^{(t)}_{b}+\mathbf{z}^{(t)}_{I,b}+\overline{\mathbf{I}}^{(t-1)}_{b}-\mathbf{z}^{(t-1)}_{I,b}\bigr)\Bigr$]
$\qquad\in\nabla f(\mathcal{X}^{(t)})+\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\boldsymbol{\psi}^{(t)}_{l,b}+N_{\mathcal{C}}(\mathcal{X}^{(t)}),$ (39a)
@@
Comparing (V-A1) and (37a), the stopping criteria for inner-level algorithm 2 will be according to (36a).
@@ V-B Convergence
To establish the convergence of the proposed two-level distributed PFBWD algorithm, we verify that the reformulated problem (IV-A) satisfies the conditions required for the convergence results in [33]. Specifically, we show that the three assumptions establishing the convergence of the two-level distributed algorithm are satisfied for the problem in this work, i.e., problem (IV-A).
Assumption 1.
Proof.
The objective function $f(\mathcal{X})=-\sum_{u\in\mathcal{U}}\log(t_{u})$ is continuously differentiable for $t_{u}>0,~\forall u$. Since the constraints of the problem explicitly enforce $t_{u}>0,~\forall u$, $f(\mathcal{X})$ is continuously differentiable over the entire feasible set. Next, consider the constraint sets $\mathcal{C}$ and $\overline{\mathcal{C}}$, defined in (21a). The set $\overline{\mathcal{C}}$ is convex because it is described by SOC constraints in (II-B) and (19a), which define convex feasible regions. Moreover, these constraints impose explicit upper and lower bounds on the optimization variables $\bar{\mathcal{X}}$, implying that $\overline{\mathcal{C}}$ is closed and bounded, and hence compact. Similarly, the set $\mathcal{C}$ is compact since the constraints (14a) and (15a) impose finite upper and lower bounds on all global optimization variables $\mathcal{X}$. As a result, $\mathcal{C}$ is closed and bounded, and therefore compact. ∎
@@
Proof.
Under Assumption 1, the constraint sets $\mathcal{C}$ and $\overline{\mathcal{C}}$ are compact and nonempty. Since problem (IV-A) only enforces the membership constraints $\mathcal{X}\in\mathcal{C}$ and $\overline{\mathcal{X}}\in\overline{\mathcal{C}}$, together with the linear coupling constraints (23a), feasibility reduces to the existence of at least one point satisfying these sets simultaneously. Because the original problem (III) admits a feasible solution, there exists at least one pair $(\mathcal{X},\overline{\mathcal{X}})$ satisfying the defining constraints of $\mathcal{C}$ and $\overline{\mathcal{C}}$. Moreover, the linear constraints (23a) do not restrict feasibility, as the auxiliary variables $\mathbf{z}_{l,b},~\forall l,~\forall b,$ are unconstrained in problem (IV-A) due to the relaxation of (24) via the augmented Lagrangian formulation. Therefore, the feasible set of problem (IV-A) is nonempty, and problem (IV-A) is feasible.
Under Assumption 1, the feasible set of problem (III) is nonempty, compact, and closed. Moreover, the objective function $f(\mathcal{X})$ is continuous and continuously differentiable over the feasible region. By the Weierstrass extreme value theorem, problem (III) admits at least one global minimizer over the feasible region. Any global minimizer is, in particular, a local minimizer. Since all constraints of problem (III) are continuously differentiable and standard constraint qualifications hold for the feasible region defined by $\mathcal{C}$ and $\overline{\mathcal{C}}$, the KKT conditions are necessary optimality conditions at any local minimizer. Therefore, there exists at least one feasible point satisfying the KKT conditions of problem (III). By definition, such a point is a stationary point. Hence, the set of stationary points of problem (III) is nonempty. ∎
@@
Proof.
At inner-level iteration $t$, the inner-level ADMM block 1 Algorithm 1 solves a nonconvex problem through SCA. By construction, the solution of Algorithm 1 is obtained by solving a convex surrogate problem generated via SCA. Each surrogate problem is a tight local approximation of the original nonconvex subproblem (30b) and shares the same first-order behavior at the previous iterate $\mathcal{X}^{(t-1)}$. Since the surrogate problem is convex and feasible, its optimal solution $\mathcal{X}^{(t)}$ satisfies the first-order optimality conditions. Moreover, due to the consistency and majorization properties of the SCA framework [30], satisfaction of the surrogate optimality conditions implies stationarity of $\mathcal{X}^{(t)}$ with respect to the original augmented Lagrangian, i.e.,
$0\in\partial_{x}L(\mathcal{X}^{(t)},\overline{\mathcal{X}}^{(t-1)},\mathcal{Z}^{(t-1)},\boldsymbol{\Psi}^{(t-1)}).$
Finally, since the augmented Lagrangian is bounded from below over the compact feasible set (following Assumption 1), the sequence of objective values is finite. Therefore, at each iteration $t$, the first-block update produces a stationary solution with a non-increasing augmented Lagrangian value.
@@ V-C Computational Complexity and Signaling Overhead
VI Numerical Results and Discussion
@@
HAPS altitude 2020 Km [43]
Center frequency ($f_{c}$) $2.545$ GHz [43]
$\sigma_{\xi}$$K_{u}$ 881010 [35]
$d_{x},~d_{y}$ $\lambda/2$
$\sigma^{2}_{n}$ 100-100 dBm
$P_{b}^{\text{max}},~\forall b\in\mathcal{B}\setminus\{B+1\},~P_{B+1}^{\text{max}}$ $43$ dBm, $52$ dBm [43]
@@
$\delta,~\omega,~\gamma$ 2,0.5,1.52,~0.5,~1.5 [33]
@@
In this section, we evaluate the performance of the proposed distributed PFBWD algorithm in HAPS-empowered vHetNets and compare it against both standalone terrestrial networks and the centralized approach. The standalone terrestrial network refers to a scenario where no HAPS is involved. Integrating HAPS into terrestrial networks under a harmonized spectrum introduces strong inter-tier interference, which raises concerns regarding potential performance degradation, compared to the standalone terrestrial networks. Therefore, it is essential to benchmark the proposed system model against a standalone terrestrial network. Moreover, comparisons with the centralized algorithm are necessary since it serves as an upper performance bound, enabling a clear assessment of the optimality loss incurred by distributed implementation. In the case of a centralized algorithm, the original problem is solved using the SCA framework, which incorporates some RLT, as detailed in [4]. The simulated network consists of four MBSs and one HAPS, deployed over a $4~\text{km}\times 4~\text{km}$ square urban area. The HAPS and MBSs are equipped with $8\times 8$ and $4\times 4$ MIMO antenna arrays, respectively, and jointly serve $U$ single-antenna UEs in a cell-free architecture. The UEs are uniformly distributed across the considered area. The remaining simulation parameters and their values are summarized in Table I. All results are averaged over $1000$ independent network realizations. The algorithm is implemented using CVX parser, which runs on MATLAB, with mosek 9.1.9 as an internal solver [41].
@@
First, we discuss the convergence behavior of the proposed distributed PFBWD algorithm. Fig. 3 plots the average sum of residuals, defined as $\sum_{b\in\mathcal{B}}\sum_{l\in\mathcal{L}}\left\|\mathbf{r}_{l,b}\right\|$, over the inner-level ADMM Algorithm 2 iterations $t$, and the average PF objective function value (i.e., $\sum_{u\in\mathcal{U}}{\log(\log_{2}(1+\gamma_{u}))}$) over the outer-level Algorithm 3 iterations $k$. It can be observed that the residual values converge to zero as $t$ increases, while the objective value increases with $k$, demonstrating the convergence behavior of the proposed algorithm. The outer-level ALM algorithm also exhibits fast convergence as the objective value saturates after a few iterations (on average $k=4$ iterations). However, the choice of penalty parameters plays a critical role in determining the number of required iterations. As shown in Table II, as the ratio of the inner-level to outer-level penalty parameters (i.e., $\delta$ in Algorithm 2) increases, the number of required iterations reduces. This is because a higher penalty parameter enforces the satisfaction of the coupling constraints more aggressively, thereby accelerating convergence to feasibility. However, excessively large values of $\delta$ may lead to numerical instability [25].
(a) CDF of SE per UE (b/s/Hz/UE).
(b) CDF of PF objective function value.
Figure 4: Statistical behavior of SE per UE and PF objective function value in a HAPS-empowered vHetNet (44 MBSs + 11 HAPS) serving $16$ UEs.
@@
TABLE III: Comparison of computational complexity and signaling overhead for centralized vs. distributed PFBWD algorithms in a vHetNet (44 MBSs + 11 HAPS) serving $16$ UEs.
@@
Figure 5: CDF of relative LPEI (proposed distributed PFBWD algorithm 3 over centralized algorithm).
Table III summarizes the computational complexity and signaling overhead requirements of the centralized and distributed algorithms when serving $16$ UEs, considering $4\times 4$ and $8\times 8$ antenna arrays at the MBSs and the HAPS, respectively. It is evident that the proposed distributed PFBWD Algorithm 3 incurs substantially lower computational complexity and significantly reduced data exchange, resulting in a pronounced reduction in signaling overhead compared to the centralized benchmark. Furthermore, Figs. 4(a) and 4(b) depict the cumulative distribution functions (CDFs) of SE per UE and the PF objective value, respectively, for a vHetNet comprising $4$ MBSs and $1$ HAPS serving $16$ UEs. As expected, the centralized algorithm achieves higher performance due to its access to global channel state information and centralized optimization. In contrast, the distributed algorithm relies solely on local CSI and limited coordination among BSs. Quantitatively, Fig. 4(a) shows that the average SE per UE achieved by the centralized algorithm is approximately $20.9$ b/s/Hz/UE, whereas the distributed algorithm attains $18.1$ b/s/Hz/UE, corresponding to a performance gap of about $13\%$. Similarly, Fig. 4(b) indicates average PF objective values of $70.5$ and $66.9$ for the centralized and distributed schemes, respectively, resulting in a gap of approximately $5\%$. The similar slopes and spreads of the CDFs indicate comparable variability across network realizations, confirming that the distributed algorithm operates close to the centralized benchmark in terms of achievable performance.
To jointly evaluate the achieved performance in terms of the objective function value, computational cost, and the signaling overhead, we define a logarithmic performance–efficiency index (LPEI) as
$\small\text{LPEI}=\frac{\text{Average Objective Function Value}}{\log_{10}(\text{Computational Complexity}\times\text{Signaling Overhead})}.$ (43)
Logarithmic normalization is applied to ensure meaningful comparison when costs between the distributed and centralized algorithms differ significantly. Accordingly, Fig. 5 plots the CDF of the relative LPEI, defined as the ratio between the LPEI of the proposed distributed PFBWD algorithm and that of the centralized one. As observed, the relative LPEI remains consistently greater than one across all network realizations, indicating that the proposed distributed PFBWD algorithm is uniformly more efficient than the centralized solution. These results confirm that, despite a modest reduction in SE and PF performance, the proposed distributed PFBWD algorithm achieves higher overall efficiency by substantially reducing computational complexity and signaling overhead, making it a scalable and practical solution for large-scale HAPS-empowered vHetNets.
(a) Average SE per UE vs. number of UEs
(b) Average PF objective function value vs. number of UEs
Figure 6: Performance comparison of centralized and distributed PFBWD algorithms for different numbers of UEs in a HAPS-empowered vHetNet (44 MBSs + 11 HAPS).
Furthermore, Fig. 6 illustrates the performance gap between the centralized and distributed PFBWD algorithms with increasing number of UEs. Specifically, Figs. 6(a) and  6(b) plot the average SE per UE and the average PF objective function value, respectively, in a HAPS-empowered vHetNet consisting of $4$ MBSs and $1$ HAPS. As the number of UEs grows, the average SE per UE decreases mainly due to two reasons: i) the available resources are shared among more UEs and ii) the increased interference. Additionally, the performance gap between the centralized and distributed algorithms slightly widens with increasing UEs. This is because the amount of missing information at each BS in the distributed approach grows, leading to a minor decline in performance. Importantly, this minor decrease in performance comes with significant gains in algorithmic complexity and reduced signaling overhead among BSs.
@@
Figure 8: Impact of HAPS antenna architecture on the average SE per UE in a vHetNet (44 MBSs + 11 HAPS) serving $16$ UEs.
Figure 9: Impact of ($P^{\text{max}}_{B+1}$) on the minimum SE per UE in a vHetNet (44 MBSs + 11 HAPS) serving $16$ UEs.
@@
In addition to the network architecture and optimization framework, the system-level parameters of the HAPS play a crucial role in determining the achievable SE in HAPS-empowered vHetNets. In particular, the antenna configuration and maximum available transmit power at the HAPS directly influence the beamforming capability, interference management effectiveness, and overall performance. To this end, Fig. 8 presents the average SE per UE in a $(4+1)$ vHetNet configuration serving $16$ UEs, for various HAPS antenna configurations. As shown, the SE gain provided by HAPS integration can be further amplified by increasing the number of antenna elements at the HAPS. This is particularly promising, as the large surface area of HAPS platforms allows for accommodating a significantly larger number of antennas, enabling improved beamforming capabilities and higher spatial resolution. Furthermore, Fig. 9 illustrates the minimum SE per UE for different maximum available transmit power, i.e., $P^{\text{max}}_{B+1}$, levels at the HAPS. The results indicate that increasing the HAPS transmit power leads to improvement in the achieved minimum SE, highlighting the importance of power provisioning at the HAPS for enhancing fairness and ensuring satisfactory service to UEs.